# Towards Adapting Open-Source Large Language Models for Expert-Level Clinical Note Generation

### v5

@@ 5.3 Analysis of RLHF
We gathered quantitative feedback from physician reviewers, who noted that the most frequent issues in the initial round of notes were inaccurate information or hallucinations. Two out of three reviewers observed that the quality of notes significantly improved with fewer hallucinations after one round DPO, describing notes as “often indistinguishable.” Consequently, we conducted only two rounds of DPO, mindful of its time-intensive nature. Appendix Figure 9 presents a qualitative analysis of a specific case at different stages of model training. This example illustrates that RLHF resulted in more granular changes, building on the results of RLAIF while maintaining the same output structures. We named the model after our RLHF step as LLaMA-Clinic.
@@ 5.4 Analysis of Physician Reader Study
We presented the results from the physician reader study in Figure 4. Gemini Pro achieved the highest scores across all three criteria (Figure 4B). Remarkably, the majority (90.4%) of the individual evaluations rated the notes generated by LLaMA-Clinic as “acceptable” or higher across all three criteria (Figure 4C). Furthermore, the overall distribution of scores was also similar among the three groups (Figure 4C). This aligns with qualitative feedback from two out of three reviewers stating that the quality of notes was, for the “majority” of the time, indistinguishable among the groups.
The metric of real-world readiness yielded intriguing observations. Physicians were asked to evaluate the notes as though they would be utilized in everyday clinical practice, assuming the physicians would proofread and make necessary edits. Compared to notes authored by physicians, LLaMA-Clinic received higher overall readiness scores, especially in the more challenging “Assessment and Plan” section (4.1 vs. 4.0), although it scored slightly lower in accuracy (Figure 4b). This finding suggest that physicians might be more tolerant of minor factual inaccuracies in a real-world setting if the notes require fewer edits.
@@ 6 Discussion
Our methodology to adapt LLM for clinical note generation revealed several key considerations. We initiated this process by establishing a “best practice” note format informed by a consensus among our physician authors. This approach tackles the variability in the style, format, and quality of physician notes, which could otherwise compromise the training of LLMs. Instead of relying on the LLM to identify the most effective note structure, we advocate for healthcare providers to set these standards themselves. It is important to acknowledge that the “best practice” note format employed in this study reflects the consensus of a group of internal medicine physicians within a single organization. This format may not be generalizable to other specialties or other organizations, such as orthopedic surgery. Furthermore, we examined a workflow in which providers critically review and refine AI-generated notes prior to their filing. Under such a workflow, physician preferences may shift toward notes requiring fewer revisions and edits, with a higher tolerance for minor inaccuracies.
@@ 7 Conclusions
Our research underscores the potential of training an open-source LLM for outpatient note generation with promising clinical applicability. Healthcare institutions are in a privileged position to undertake such endeavors, given their access to extensive EMR data and a wealth of domain expertise critical for implementing RLHF. Our work was based on fewer than 1,500 patient-doctor dialogues and limited physician preference data. When implementing a similar project in a healthcare institution, these training data could be scaled up by several orders of magnitude, likely boosting performance further. Furthermore, the prospect of conducting similar work for other clinical note-generation tasks, such as creating discharge summaries for hospitalized patients, is particularly exciting.
8 Code and Data Availability
Code and data for the current study are available at https://anonymous.4open.science/r/llama-clinic-CE31.
@@ 9 Ethical Considerations
We followed the same approach as in van2023clinical; walsh2017measuring to assess the potential consequences of factual errors. One physician evaluated the likelihood and severity of harm associated with notes that received an accuracy or completeness score below 5 in his review (n = 22, 31 and 39 for Gemini Pro, LLaMA-Clinic, and Physician, respectively). All notes within the LLaMA-clinic and physician groups were assigned scores of “None” for the extent of harm and “Low” for the likelihood of harm. In contrast, the Gemini Pro group contained a single case rated with “Mild to Moderate” extent of harm and a “Medium” likelihood of harm.
@@ 10 Limitations
Our study serves as a proof of concept and encounters limitations, notably the scarcity of publicly available patient-doctor dialogues for model training, with the largest ACI-BENCH dataset comprising fewer than 300 cases. Additionally, the effectiveness of outpatient note generation is fundamentally linked to the content of patient-doctor dialogues. Although the ACI-BENCH data may have synthesized dialogues to encapsulate all necessary information for composing a comprehensive note, such ideal conditions may not always reflect real-world scenarios. For instance, time constraints may prevent physicians from discussing every detail of medical reasoning with patients, potentially degrading the quality of the generated notes due to the suboptimal input dialogue. Our research was also constrained by limited availability of physician evaluators, which restricted our ability to conduct extensive hyperparameter searches or additional rounds of RLHF. Our final evaluation was conducted with only three physicians due to the task’s time-intensive nature.
@@ A.2.2 CP Experiment 2
Approach: In this experiment, we implemented mixed precision training, wherein weights and gradients were stored in bf16 format, and optimizer states in fp32. This decision was based on the hypothesis that training with pure bf16 might pose challenges in achieving convergence zamirai2020revisiting. The remaining setup parameters were consistent with those outlined in Experiment 1.
@@ A.2.5 CP Experiment 5
Approach: Motivated by the inferior performance of LoRA compared to full parameter training across various benchmarks han2023medalpaca, we explored the implementation of full parameter training on 7B models, considering their similar VRAM requirements (compared to LoRA training on the 13B models).
@@ A.2.6 CP Experiment 6
Approach: We reduced LR to 2e-5, down from 3e-4 in the previous experiment. Of note, the vanilla LLaMA-2 7B model’s used a LR of 3e-4 for pretraining touvron2023llama.
@@ A.3.1 SFT Experiment 1
    7B_short 0.2332 0.3373 0.3163 0.2730 0.2259 0.2209 0.3164 0.3410
    7B_long 0.2308 0.3619 0.0875 0.3262 0.2030 0.1992 0.2357 0.3205
    7B_chat_short 0.2644 0.3471 0.0292 0.4465 0.2531 0.2670 0.3370 0.4436
    7B_chat_long 0.2322 0.4270 0.4072 0.2639 0.2305 0.2841 0.3193 0.3916
    13B_short 0.1335 0.3472 0.2353 0.3757 0.1586 0.1873 0.0540 0.3079
    13B_long 0.1929 0.3329 0.2234 0.2949 0.1816 0.2602 0.2367 0.2918
    13B_chat_short 0.2161 0.3703 0.4168 0.2297 0.2411 0.3901 0.3629 0.3202
    13B_chat_long 0.1797 0.3617 0.3342 0.4065 0.2337 0.3381 0.3944 0.3903
@@ A.3.2 RLAIF Experiment 1
Results: Performance metrics following RLAIF Experiment 1 are presented in Appendix Table 2. Although we achieved commendable performance by the third round of DistillDirect (notably the best performance came from the 7B-chat_short model), the training process exhibited instability. For instance, the performance in the second round of DistillDirect for the 7B_chat_short model showed a significant decline from the previous round. Across all models, we did not observe a consistent enhancement in performance (as assessed by the ROUGE-1 score on the validation set) with additional rounds of DPO training. Notably, training accuracy reached 100% prematurely, during the first epoch for all models with a high reward margin (as shown in Appendix Figure 6), indicating potential overfitting.
@@
Appendix Figure 6: Example Training Set Accuracy and Reward Margin during DPO with a LR of 2e-5. Examples taken from 13B_chat_short_R3, and all other runs have similar training curve with high accuracy and reward margin early on with this LR.
@@ A.3.3 SFT Experiment 2
Approach: Through discussions with our physician co-authors, we concluded that before AI can assist physicians in generating notes, the physicians themselves must determine what constitutes a “best practice” note format. It’s acknowledged that the “best practice” might diverge from existing notes for valid reasons (notably, recognizing that physicians can also produce suboptimal notes). With this perspective, we revised our prompts and provided a single-shot example to Gemini-pro to foster “constrained” note generation tailored towards a specific style. This new note style was collectively endorsed by our physician authors as superior to many examples within ACI-BENCH. Consequently, we directed Gemini-pro to generate new reference notes from the dialogues in ACI-BENCH, resulting in a dataset we named Modified ACI-BENCH and discontinuing the use of its original notes.. At this stage, we also regenerated reference notes for Dialogue-G using Gemini Pro. Initially, we considered having our physicians edit Gemini-pro’s notes before employing them for training. However, we abandoned this idea due to the time-intensive nature of the task and the satisfactorily high quality of outputs from Gemini-pro with our revised prompt.
@@
 7B_short 0.1998 0.4894 0.5198 0.3792 0.2617 0.4885 0.0917 0.4100
 7B_long 0.1949 0.4688 0.0594 0.3920 0.2446 0.3705 0.3053 0.4040
 7B_chat_short 0.2961 0.4816 0.4946 0.0586 0.3157 0.5201 0.4763 0.3867
 7B_chat_long 0.3067 0.5051 0.0964 0.0773 0.2988 0.4907 0.4275 0.3628
 13B_short 0.2463 0.5064 0.4904 0.4066 0.2277 0.5252 0.4059 0.4493
 13B_long 0.2525 0.4199 0.5113 0.3309 0.2565 0.4840 0.4610 0.3566
 13B_chat_short 0.3475 0.5115 0.5030 0.5441 0.3055 0.5277 0.4422 0.5422
 13B_chat_long 0.3463 0.4856 0.5308 0.3863 0.3620 0.4752 0.4730 0.4633
@@ A.3.4 RLAIF Experiment 2
Results: A number of intriguing observations can be made from this iteration (Appendix Table 3).
@@
  • 7B vs 13B models: With the current LR, the 7B models exhibited significant instability during DistillDirect. For instance, the 7B_short model demonstrated a notable performance degradation from round 1 to round 2. Upon manual evaluation of its outputs, the model incorrectly generated “Subjective” sections when tasked with producing “Assessment and Plan” and frequently returned null outputs. Given the more stable training observed with 13B models utilizing LoRA, along with their superior performance and reduced storage requirements compared to full parameter training for 7B models, we have decided to proceed exclusively with 13B models henceforth.
@@
  • Best Performer by ROUGE-1 : Among our CP models, the 13B_chat model pre-trained with the Discharge-short dataset emerged as the top performer. However, the overall best model was the 13B_chat model following SFT and three rounds of DistillDirect, but without CP. It achieved ROUGE-1 scores of 0.5452 for “Subjective” and 0.5597 for “Assessment and Plan”. Naturally, this led us to question the value of CP in our specific task. Such skepticism seemed justified, particularly given the differences between our training corpus of discharge summaries and that of outpatient clinical notes, which could be deemed out-of-distribution for our task. Interestingly, a more thorough review and manual analysis of the notes by our physician author unveiled intriguing subtleties.
@@
  • Possible Best Performer by Physician Review: The quality of the notes generated by the top-performing models (e.g., 13B_chat_R3, 13B_short_R1, 13B_chat_short_R3, 13B_chat_long_R2) appears nearly indistinguishable to our physician author. Overall, we find the “Subjective” sections of these notes to be satisfactory and potentially suitable for clinical application. However, the primary deficiency in achieving clinical-ready notes lies within the “Assessment and Plan” sections, particularly regarding medical reasoning and certain linguistic nuances. Although somewhat subjective, we felt the outputs from 13B_short_r1 bear the greatest resemblance to an actual provider’s note. This finding underscores the limitations of quantitative metrics in tasks involving domain-specific language generation.
@@
  • Persistent DistillDirect Training Instability: Despite the creation of a more uniform dataset, DistillDirect training remained susceptible to fluctuations. This instability was more pronounced for 7-billion parameter models but was also observed in 13-billion parameter models (Appendix Table 3). The 13B_long model serves as an example, where the ROGUE-1 score dropped from 0.5113 in round 2 to 0.3309 in round 3 for the “Subjective” section. Reducing the number of epochs per DistillDirect round from three to one did not mitigate this issue. In all DistillDirect rounds, training set accuracy rapidly reached a perfect score (1.0) halfway through the first epoch, accompanied by a high reward margin (as illustrated in the initial third of the curve shown in Appendix Figure 6). This observation again suggests potential overfitting. Interestingly, similar rapid accuracy increases were reported during Zephyr training, and overfitting did not negatively impact performance on downstream tasks tunstall2023zephyr.
@@
  • Reflection on LR Selection: The optimal LR for DPO has yet to be established in literature, with only a limited number of LLMs having undergone DPO training to date. Notably, given the small size of our training set, our initial LR choice of 2e-5 is larger than those selected in Zephyr and Tulu-2, where both models opted for an LR of 5e-7 ivison2023camels; tunstall2023zephyr. The Tulu-2 paper mentioned that a “slow LR… is required for stable and effective DPO training.” Conversely, in another study, Gaudi-2 employed a significantly higher LR (5e-4, compared to an LR of 1e-4 during its SFT), despite using the same training data as Zephyr, and still achieved a favorable response intel_analytics_2023.
@@
Solution: Ideally, at this stage, our goal would be to select the best model, sample its generations under various settings, and forward these samples to physicians for labeling in preparation for the RLHF. However, the absence of a stable DistillDirect setup presents a significant challenge. Moreover, there is no assurance that further DPO training with real human preference data will result in performance improvement. Additionally, we have yet to confirm that we have identified the optimal model using the existing training dataset. Therefore, we have decided to conduct a limited LR search specifically for DistillDirect.
@@
 13B_short 0.4006 0.5423 0.5186 0.4702 0.4032 0.5683 0.4039 0.5242
 13B_long 0.3570 0.4799 0.5779 0.3981 0.4298 0.5266 0.4688 0.3780
 13B_chat_short 0.4390 0.5455 0.5248 0.5560 0.4825 0.5565 0.5236 0.5694
 13B_chat_long 0.4272 0.5149 0.5423 0.4188 0.4459 0.5487 0.5322 0.4852
@@
 13B_chat 0.4077 0.5351 0.5582 0.5466 0.4383 0.5522 0.4907 0.5847
Appendix Table 4: ROUGE-1 Score after RLAIF Experiment 2 with Lower Temperature. Performance reported on validation subset of modified ACI-BENCH. R0 represent the model after SFT, with or without CP. R1 to R3 represents performance after respective rounds of DistillDirect. Models ended in long were pretrained using Discharge-long dataset, while models ended in short were pretrained using Discharge-short dataset. Bolded scores denote the best performance with respect to the task. The experimental setup includes training on the new reference notes (modified ACI-BENCH) with a LR of 2e-5, and 1 epoch per training round. The temperature is set at 0.6 during generation time to calculate ROUGE-1, using the same model checkpoints as in Appendix Table 3

### v6

@@ 5.3 Analysis of RLHF
We collected quantitative feedback from physician reviewers, who identified inaccurate information and hallucinations as the most common issues in the LLM-generated notes. Two out of three reviewers noted a significant improvement in the quality of the notes, with fewer hallucinations after one round of DPO. Consequently, we conducted only two rounds of DPO, mindful of its time-intensive nature (Appendix Figure 6A). We present a qualitative analysis of a specific case across different stages of model training, as shown in Appendix Figure 5. This example illustrates that RLHF introduced more granular changes, building on the outcomes of RLAIF while preserving the original style and structure. We named the model after our RLHF step as LLaMA-Clinic.
@@ 5.4 Analysis of Physician Reader Study
We presented the results from the physician reader study in Figure 4. Gemini Pro achieved the highest scores across all three criteria (Figure 4B). Remarkably, the majority (92.8%) of the individual evaluations rated the notes generated by LLaMA-Clinic as “acceptable” or higher across all three criteria (Figure 4C). Furthermore, the overall distribution of scores was also similar among the three groups. This aligns with qualitative feedback from three out of four reviewers, who stated that the quality of notes was, for the majority of the time, indistinguishable among the groups, particularly in the “Subjective” section.
The metric of real-world readiness yielded intriguing observations. Physicians were asked to evaluate the notes as though they would be utilized in everyday clinical practice, assuming the physicians would proofread and make necessary edits. LLaMA-Clinic achieved comparable overall readiness scores to physician-authored notes, including in the more complex “Assessment and Plan” section, as shown in Figure 4B. Overall, LLaMA-Clinic received higher completeness scores but lower accuracy scores compared to physician-authored notes. This finding suggests that physicians might be more tolerant of minor factual inaccuracies in a real-world setting if the notes require fewer edits.
@@ 6 Discussion
Our methodology for adapting LLMs to clinical note generation revealed several key considerations. We initiated this process by establishing a “best practice” note format informed by a consensus among our physician authors. This approach tackles the variability in the style, format, and quality of physician notes, which could otherwise compromise the training of LLMs. Instead of relying on the LLM to identify the most effective note structure, we advocate for healthcare providers to set these standards themselves. It is important to acknowledge that the “best practice” note format employed in this study reflects the consensus of a group of internal medicine physicians within a single organization. This format may not be generalizable to other specialties or other organizations, such as orthopedic surgery. Furthermore, we examined a workflow in which providers critically review and refine AI-generated notes prior to their filing. Under such a workflow, physician preferences may shift toward notes requiring fewer revisions and edits, with a higher tolerance for minor inaccuracies.
@@ 7 Conclusions
Our research highlights the potential of training an open-source LLM for outpatient note generation, showing strong promise for real-world clinical applications. Healthcare institutions are in a privileged position to undertake such endeavors, given their access to extensive EMR data and a wealth of domain expertise critical for implementing RLHF. Our work was based on fewer than 1,500 patient-doctor dialogues and limited physician preference data. When implementing a similar project in a healthcare institution, the training data could be scaled up significantly, potentially leading to further performance improvements. Lastly, the prospect of extending this work to other clinical note-generation tasks, such as creating discharge summaries for hospitalized patients, is particularly exciting.
8 Acknowledgments
This research was supported by NSF awards SCH-2205289. The funder played no role in the study design, data collection, analysis, and interpretation of data, or the writing of this manuscript. This research is also supported by the Protected Research Time Grant from the Southwest Minnesota Region of the Mayo Clinic Health System.
@@ 9 Ethical Considerations
We followed the same approach as in Van Veen et al. (2023); Walsh et al. (2017) to assess the potential consequences of factual errors. One physician evaluated the likelihood and severity of harm associated with notes that received an accuracy or completeness score below 5 in his review (n = 22, 31 and 39 for Gemini Pro, LLaMA-Clinic, and Physician, respectively). All notes within the LLaMA-clinic and physician groups were assigned scores of “None” for the extent of harm and “Low” for the likelihood of harm. In contrast, the Gemini Pro group contained a single case rated with “Mild to Moderate” extent of harm and a “Medium” likelihood of harm.
@@ 10 Limitations
Our study serves as a proof of concept and encounters limitations, notably the scarcity of publicly available patient-doctor dialogues for model training, with the largest ACI-BENCH dataset comprising fewer than 300 cases. Additionally, the effectiveness of outpatient note generation is fundamentally linked to the content of patient-doctor dialogues. Although the ACI-BENCH data may have synthesized dialogues to encapsulate all necessary information for composing a comprehensive note, such ideal conditions may not always reflect real-world scenarios. For instance, time constraints may prevent physicians from discussing every detail of medical reasoning with patients, potentially degrading the quality of the generated notes due to the suboptimal input dialogue. Our research was also constrained by limited availability of physician evaluators, which restricted our ability to conduct extensive hyperparameter searches or additional rounds of RLHF. Our final evaluation was conducted with only four physicians due to the task’s time-intensive nature.
@@ A.2.2 CP Experiment 2
Approach: In this experiment, we implemented mixed precision training, wherein weights and gradients were stored in bf16 format, and optimizer states in fp32. This decision was based on the hypothesis that training with pure bf16 might pose challenges in achieving convergence Zamirai et al. (2020). The remaining setup parameters were consistent with those outlined in Experiment 1.
@@ A.2.5 CP Experiment 5
Approach: Motivated by the inferior performance of LoRA compared to full parameter training across various benchmarks Han et al. (2023), we explored the implementation of full parameter training on 7B models, considering their similar VRAM requirements (compared to LoRA training on the 13B models).
@@ A.2.6 CP Experiment 6
Approach: We reduced LR to 2e-5, down from 3e-4 in the previous experiment. Of note, the vanilla LLaMA-2 7B model’s used a LR of 3e-4 for pretraining Touvron et al. (2023).
@@ A.3.1 SFT Experiment 1
    7B-short 0.2332 0.3373 0.3163 0.2730 0.2259 0.2209 0.3164 0.3410
    7B-long 0.2308 0.3619 0.0875 0.3262 0.2030 0.1992 0.2357 0.3205
    7B-chat-short 0.2644 0.3471 0.0292 0.4465 0.2531 0.2670 0.3370 0.4436
    7B-chat-long 0.2322 0.4270 0.4072 0.2639 0.2305 0.2841 0.3193 0.3916
    13B-short 0.1335 0.3472 0.2353 0.3757 0.1586 0.1873 0.0540 0.3079
    13B-long 0.1929 0.3329 0.2234 0.2949 0.1816 0.2602 0.2367 0.2918
    13B-chat-short 0.2161 0.3703 0.4168 0.2297 0.2411 0.3901 0.3629 0.3202
    13B-chat-long 0.1797 0.3617 0.3342 0.4065 0.2337 0.3381 0.3944 0.3903
@@ A.3.2 RLAIF Experiment 1
Results: Performance metrics following RLAIF Experiment 1 are presented in Appendix Table 1. Although we achieved commendable performance by the third round of DistillDirect (notably the best performance came from the 7B-chat_short model), the training process exhibited instability. For instance, the performance in the second round of DistillDirect for the 7B-chat-short model showed a significant decline from the previous round. Across all models, we did not observe a consistent enhancement in performance (as assessed by the ROUGE-1 score on the validation set) with additional rounds of DPO training. Notably, training accuracy reached 100% prematurely, during the first epoch for all models with a high reward margin (as shown in Appendix Figure 2), indicating potential overfitting.
@@
Appendix Figure 2: Example Training Set Accuracy and Reward Margin during DPO with a LR of 2e-5. Examples taken from 13B-chat-short_R3, and all other runs have similar training curve with high accuracy and reward margin early on with this LR.
@@ A.3.3 SFT Experiment 2
Approach: Through discussions with our physician co-authors, we concluded that before AI can assist physicians in generating notes, the physicians themselves must determine what constitutes a “best practice” note format. It’s acknowledged that the “best practice” might diverge from existing notes for valid reasons (notably, recognizing that physicians can also produce suboptimal notes). With this perspective, we revised our prompts and provided a single-shot example to Gemini-pro to foster “constrained” note generation tailored towards a specific style. This new note style was collectively endorsed by our physician authors as superior to many examples within ACI-BENCH. Consequently, we directed Gemini-pro to generate new reference notes from the dialogues in ACI-BENCH, resulting in a dataset we named Modified ACI-BENCH. At this stage, we also regenerated reference notes for Dialogue-G using Gemini Pro. Initially, we considered having our physicians edit Gemini-pro’s notes before employing them for training. However, we abandoned this idea due to the time-intensive nature of the task and the satisfactorily high quality of outputs from Gemini-pro with our revised prompt.
@@
 7B-short 0.1998 0.4894 0.5198 0.3792 0.2617 0.4885 0.0917 0.4100
 7B-long 0.1949 0.4688 0.0594 0.3920 0.2446 0.3705 0.3053 0.4040
 7B-chat-short 0.2961 0.4816 0.4946 0.0586 0.3157 0.5201 0.4763 0.3867
 7B-chat-long 0.3067 0.5051 0.0964 0.0773 0.2988 0.4907 0.4275 0.3628
 13B-short 0.2463 0.5064 0.4904 0.4066 0.2277 0.5252 0.4059 0.4493
 13B-long 0.2525 0.4199 0.5113 0.3309 0.2565 0.4840 0.4610 0.3566
 13B-chat-short 0.3475 0.5115 0.5030 0.5441 0.3055 0.5277 0.4422 0.5422
 13B-chat-long 0.3463 0.4856 0.5308 0.3863 0.3620 0.4752 0.4730 0.4633
@@ A.3.4 RLAIF Experiment 2
Results: Several intriguing observations emerge from this iteration (Appendix Table 2).
@@
  • 7B vs 13B models: With the current LR, the 7B models exhibited significant instability during DistillDirect. For instance, the 7B-short model demonstrated a notable performance degradation from round 1 to round 2. Upon manual evaluation of its outputs, the model incorrectly generated “Subjective” sections when tasked with producing “Assessment and Plan” and frequently returned null outputs. Given the more stable training observed with 13B models utilizing LoRA, along with their superior performance and reduced storage requirements compared to full parameter training for 7B models, we have decided to proceed exclusively with 13B models henceforth.
@@
  • Best Performer by ROUGE-1 : Among our CP models, the 13B-chat model pre-trained with the Discharge-short dataset emerged as the top performer. However, the overall best model was the 13B-chat model following SFT and three rounds of DistillDirect, but without CP. It achieved ROUGE-1 scores of 0.5452 for “Subjective” and 0.5597 for “Assessment and Plan”. Naturally, this led us to question the value of CP in our specific task. Such skepticism seemed justified, particularly given the differences between our training corpus of discharge summaries and that of outpatient clinical notes, which could be deemed out-of-distribution for our task.
@@
  • Possible Best Performer by Physician Review: The quality of the notes generated by the top-performing models (e.g., 13B-chat_R3, 13B-short_R1, 13B-chat-short_R3, 13B-chat-long_R2) appears nearly indistinguishable to our physician author. Overall, we find the “Subjective” sections of these notes to be satisfactory and potentially suitable for clinical application. However, the primary deficiency in achieving clinical-ready notes lies within the “Assessment and Plan” sections, particularly regarding medical reasoning and certain linguistic nuances. Although somewhat subjective, we felt the outputs from 13B-short_r1 bear the greatest resemblance to an actual provider’s note. This finding underscores the limitations of quantitative metrics in tasks involving domain-specific language generation.
@@
  • Persistent DistillDirect Training Instability: Despite the creation of a more uniform dataset, DistillDirect training remained susceptible to fluctuations. This instability was more pronounced for 7-billion parameter models but was also observed in 13-billion parameter models (Appendix Table 2). The 13B-long model serves as an example, where the ROGUE-1 score dropped from 0.5113 in round 2 to 0.3309 in round 3 for the “Subjective” section. Reducing the number of epochs per DistillDirect round from three to one did not mitigate this issue. In all DistillDirect rounds, training set accuracy rapidly reached a perfect score of 1.0 halfway through the first epoch, accompanied by a high reward margin (as illustrated in the initial third of the curve shown in Appendix Figure 2). This observation again suggests potential overfitting. Interestingly, similar rapid accuracy increases were reported during Zephyr training, and overfitting did not negatively impact performance on downstream tasks Tunstall et al. (2023).
@@
  • Reflection on LR Selection: The optimal LR for DPO has yet to be established in literature, with only a limited number of LLMs having undergone DPO training to date. Notably, given the small size of our training set, our initial LR choice of 2e-5 is larger than those selected in Zephyr and Tulu-2, where both models opted for an LR of 5e-7 Ivison et al. (2023); Tunstall et al. (2023). The Tulu-2 paper mentioned that a “slow LR… is required for stable and effective DPO training.” Conversely, in another study, Gaudi-2 employed a significantly higher LR (5e-4, compared to an LR of 1e-4 during its SFT), despite using the same training data as Zephyr, and still achieved a favorable response Intel Analytics Software (2023).
@@
Solution: Conduct a limited LR search specifically for DistillDirect.
@@
 13B-short 0.4006 0.5423 0.5186 0.4702 0.4032 0.5683 0.4039 0.5242
 13B-long 0.3570 0.4799 0.5779 0.3981 0.4298 0.5266 0.4688 0.3780
 13B-chat-short 0.4390 0.5455 0.5248 0.5560 0.4825 0.5565 0.5236 0.5694
 13B-chat-long 0.4272 0.5149 0.5423 0.4188 0.4459 0.5487 0.5322 0.4852
@@
 13B-chat 0.4077 0.5351 0.5582 0.5466 0.4383 0.5522 0.4907 0.5847
Appendix Table 3: ROUGE-1 Score after RLAIF Experiment 2 with Lower Temperature. Performance reported on validation subset of modified ACI-BENCH. R0 represent the model after SFT, with or without CP. R1 to R3 represents performance after respective rounds of DistillDirect. Models ended in long were pretrained using Discharge-long dataset, while models ended in short were pretrained using Discharge-short dataset. Bolded scores denote the best performance with respect to the task. The experimental setup includes training on the new reference notes (modified ACI-BENCH) with a LR of 2e-5, and 1 epoch per training round. The temperature is set at 0.6 during generation time to calculate ROUGE-1, using the same model checkpoints as in Appendix Table 2.